OpenStudy (anonymous):

simple trig proof?? (sorry i'm so bad at these...) tan^2(x)= 1-cos(2x)/1+cos(2x)

OpenStudy (anonymous):

r.h.s = 2sin^2 (x) / 2 cos^2 (x) = 2 tan^2 (x)

OpenStudy (anonymous):

oh sry no 2

OpenStudy (anonymous):

no 2 where?

OpenStudy (anonymous):

in the answer...there should be no 2

OpenStudy (anonymous):

r.h.s = 2sin^2 (x) / 2 cos^2 (x) = tan^2 (x)